OpenStudy (anonymous):

(Use g=9.8m/s^2.) If I weigh 707N on earth and 5320N on the surface of a nearby planet,what is the acceleration due to gravity on that planet?

OpenStudy (anonymous):

\[F=ma\] mass unchanging \[\frac{ 707 }{ 9,8 }=m\] use same mass in \[5320=(m)(a)\] get a

OpenStudy (anonymous):

alternatively by propotion \[\frac{ 707 }{ 9,8 }=\frac{ 5320 }{ a }\] solve a

OpenStudy (anonymous):

Thanks much.